Calculate the value of the discriminant of x^2 + x + 3 = 0 By examining the sign of the discriminant in the previous equation, how many x-intercepts would the graph of y = x^2 +x + 3 have? why?
Answer by stanbon(75887)

Calculate the value of the discriminant of x^2 + x + 3 = 0
b^2 - 4ac = 1^2-4*1*3 = 1-12 = -11
----------------------------------------
By examining the sign of the discriminant in the previous equation, how many x-intercepts would the graph of y = x^2 +x + 3 have?
None
----------------------------------------
why?
The x-intercepts are not Real Numbers; they are Complex numbers
and they do not appear on the Real Number x-axis.
=======================
Cheers,
Stan H.